OpenStudy (anonymous):

suppose F(X)=4x-2 and G(X)=-2x+1 find the value of f(0)/G(4)

OpenStudy (anonymous):

For F(x) = 4x-2 then f(0) = 4(0) - 2 = 0-2 = -2 For g(x) = -2x +1 then g(4) = -2(4) +1 = -8 +1 = -7 therefore F(x) = -2 & g(4) = -7

OpenStudy (anonymous):

Gx1

OpenStudy (anonymous):

F(0)=-2 G(4)=-7 F(0)/G(4)=2/7